# From $U(1)\times U(1)$ Symmetry Breaking to Majoron Cosmology: Insights from NANOGrav 15-year Data

### v1

@@ Abstract
We study the cosmology of a modified majoron model motivated by the need to protect a global $U(1)$ symmetry from gravity-induced hard explicit breaking (by $d\leq 4$ operators) at the Planck scale. The model extends the Standard Model by introducing a gauged $U(1)_{B-L}$ and an approximate global $U(1)$ symmetry, each spontaneously broken by a corresponding complex scalar singlet. This setup gives rise to a network of effectively global and local cosmic strings, whose stochastic gravitational wave signals can jointly account for the spectrum observed by the NANOGrav collaboration, particularly for majoron masses $m_{\chi}<10^{-23}$ eV. Although the fit is not as strong as that from supermassive black hole mergers, the model still provides an alternative explanation rooted in high-energy physics. The model also generates light neutrino masses via the seesaw mechanism and avoids cosmological constraints from $\Delta N_{\text{eff}}$, CMB anisotropies, and isocurvature fluctuations. Although the majoron can contribute to dark matter through thermal, coherent oscillation, and string-induced production mechanisms, its relic abundance remains subdominant in the NANOGrav-compatible region. In contrast, the measured dark matter relic density is achievable at higher $m_{\chi}$, though at the cost of tension with cosmological bounds. If the NANOGrav fits are viewed as constraints, given their comparatively lower Bayes factors, they yield bounds that are significantly stronger than those imposed by the CMB and other cosmological data.
@@ I Introduction
The first direct detection of gravitational waves (GWs) by the LIGO–Virgo collaboration in 2015 [1] opened a new era in observational astronomy. In 2023, pulsar timing array (PTA) collaborations— the Parkes PTA [2], the European PTA [3], the Chinese PTA [4], and North American Nano-hertz Observatory for Gravitational Waves (NANOGrav) [5, 6]—reported evidence for a stochastic GW background in the nano-Hertz range. PTAs detect such signals by monitoring the arrival-time fluctuations of radio pulses from millisecond pulsars across the galaxy [7, 8, 9], with angular correlations following the Hellings–Downs curve [10] expected from spin-2 GWs. Among these, NANOGrav provides the most compelling evidence using 67 pulsars. While mergers of supermassive black hole binaries (SMBHBs) remain the leading astrophysical explanation [11, 12], no individual SMBHB mergers have been detected, motivating interest in cosmological sources [13, 14]. These include domain walls [15, 16], phase transitions [17, 18], scalar-induced GWs [19, 20], and cosmic strings [21, 22]. In this work, we focus on cosmic strings (CSs), which not only provide a plausible source of the signal but also offer a rare probe of high-scale physics, spanning $10^{12}$$10^{15}$ GeV (depending on local or global CS), well beyond the reach of current and proposed future colliders.
A recent work on global cosmic strings (GCSs) [23] shows that their GW spectrum, for symmetry-breaking scales around $\mathcal{O}(10^{15})$ GeV, can fall within the nano-Hertz range probed by PTAs. Intriguingly, the classic and simplest majoron model [24]—a well-known framework for neutrino mass generation—features a global $U(1)_{L}$ symmetry broken at $\mathcal{O}(10^{14}$$10^{15})$ GeV by a complex scalar $\phi$. This symmetry breaking has two key implications: (i) it generates GCSs that may emit GWs detectable by PTAs such as NANOGrav, and (ii) it induces high-scale Majorana masses for right-handed neutrinos (RHNs), enabling light neutrino mass generation via the Type-I seesaw mechanism [25, 26, 27, 28, 29, 30, 31]. The majoron is identified with the phase of the scalar field $\phi$. While the GCS-induced GW spectrum, characterised by a negative spectral tilt, provides a less favourable fit to the latest data than SMBHB mergers, it still offers a promising opportunity to explore and constrain the parameter space of the majoron model. Additionally, majoron models also have a wide range of astrophysical, cosmological, and laboratory implications [32]
@@
To address these issues, we adopt a more robust framework, which we call the modified majoron model, by gauging the $U(1)_{B-L}$ symmetry and introducing an additional complex scalar $\phi^{\prime}$ [35]. This extended model includes both a local $U(1)_{B-L}$ and an approximate global $U(1)$ symmetry. Both $\phi$ and $\phi^{\prime}$ are charged under $U(1)_{B-L}$, with only $\phi$ carrying the global charge. Assigning $(B-L)$ charge of 2 to $\phi$ enables Majorana mass generation for RHNs. The $(B-L)$ charge of $\phi^{\prime}$ can be assigned such that the leading $U(1)_{B-L}$-invariant operator that breaks the global $U(1)$ symmetry arises only at dimension $d\geq 5$. The phase of $\phi$ still constitutes the majoron in this model and is the potential DM candidate. While the primary focus of this work is on the cosmological implications of the modified majoron model, we also briefly compare the results with the same from the simplest majoron model.
@@
The extended symmetry structure of the modified majoron model gives rise to the formation of intricate topological defects in the early Universe. When the local $U(1)_{B-L}$ symmetry breaks, it generates LCSs, followed by the formation of GCSs from the spontaneous breaking of the global $U(1)$. As the Universe cools down to a temperature where the Hubble constant is comparable to the majoron mass, explicit symmetry breaking becomes relevant, leading to the formation of domain walls (DWs). Each of these topological defects can emit gravitational radiation. The string-wall network’s fate hinges on how many walls attach to each string. If stable DWs form, they can dominate the Universe’s energy density, contradicting the CMB observations. However, if only a single DW attaches to either type of string, the network becomes unstable and promptly collapses [35, 58, 59]. In this paper, we analyse scenarios where only a single DW is formed so that CSs predominantly generate the GW spectrum.
@@ II.1 The simplest majoron model
[…]
where $\tilde{\Phi}$ is the dual of the SM Higgs doublet, and $N_{I},\nu_{J}$ are 4-component Majorana spinors (i.e $N^{c}_{I}=N_{I}$ and $\nu^{c}_{J}=\nu_{J}$). The indices $I,J=1,2,3$ are the generation indices. $V_{1}(\phi)$ is the tree-level potential of the singlet scalar $\phi$. In general, the Majorana-type Yukawa coupling matrix $\lambda$ can contain off-diagonal entries. However, for simplicity, we again assume the $\lambda$ matrix to be diagonal (hereafter, $\lambda_{II}$ will be simplified to $\lambda_{I}$). In this model, the lepton number is a conserved charge since all the SM lepton doublets and RHNs possess $U(1)_{L}$ charge of $L=1$ and $\phi$ carries $L=-2$. Consequently, $\phi$ interacts with RHNs only.
@@ II.2 The modified majoron model
The reduced coupling $\tilde{g}_{d,\text{grav}}$ is exponentially suppressed by the combined scalar and gravitational Euclidean action in the region near the wormhole and is expected to be very small. For $d=5-20$ and $\lambda=1$, $\tilde{g}_{d,\text{grav}}\sim 10^{-(d+2)}$. Given the structure of the effective Lagrangian in Eq. (4), it is unclear whether a single scalar field can suppress $d\leq 4$ effective operators and thus prevent explicit hard breaking of the global symmetry at the Planck scale.
To avoid this lack of clarity, we prefer a conservative approach and choose a model where a local gauge symmetry ensures that no $d\leq 4$ effective operators are induced by gravity. Such a model was constructed in Ref. [35] by gauging $U(1)_{B-L}$ and introducing an additional complex scalar field, $\phi^{\prime}$. The charge assignment of $\phi^{\prime}$ under $(B-L)$ can be chosen such that the lowest-dimensional $U(1)_{B-L}$-invariant operator that breaks the global $U(1)$ symmetry appears only at $d\geq 5$.
In a nutshell, this model extends the SM gauge group by introducing a local $U(1)_{B-L}$ symmetry and an approximate global $U(1)$ symmetry. The particle content is extended by three RHNs and two complex scalars, $\phi$ and $\phi^{\prime}$, which are all singlets under the SM gauge groups. Here, both $\phi$ and $\phi^{\prime}$ are charged under the local $U(1)_{B-L}$ symmetry, and only $\phi$ is charged under the accidental global symmetry. If one assigns $(B-L)$ charge of $q=2$ to $\phi$, it again couples with RHNs and generates Majorana masses for them. On the other hand, we keep $q^{\prime}$, the $(B-L)$ charge of $\phi^{\prime}$, as general as possible (including fractional charges) to allow us to construct different models for a given $d$-dimensional soft global symmetry breaking operator.
The lowest dimensional quantum gravity induced global symmetry breaking, but the $(B-L)$ invariant operator is
[…]
@@
[…]
@@
Once the $U(1)_{B-L}$ charge of $\phi$ is fixed at $q=2$ from the requirement of Majorana mass generation of RHNs, one is left with the freedom of only one parameter choice at their disposal to introduce a global symmetry breaking term in dimension $d$. One can thus choose to vary the parameter $n$, which can admit positive integer values up to $d-1$, and the values of $n^{\prime}$ and $q^{\prime}$ can be updated accordingly using the dimensional and gauge invariance conditions mentioned above. Such terms in the Lagrangian will inevitably give rise to an $M_{Pl}^{d-4}$-suppressed majoron mass [35]. For example, in dimension-5 for $q=2$, the $(B-L)$ invariant but global symmetry breaking term can arise if $q^{\prime}$ takes values {1/2, 4/3, 3, 8}. In contrast, if we assign a fractional $U(1)_{B-L}$ charge to $\phi^{\prime}$, say $q^{\prime}=1/8$, then the lowest dimension at which the global symmetry can break explicitly is 17. More phenomenologically motivated benchmark points are shown in Table 4 in Section VI.
@@ III Cosmic Strings and Gravitational Waves
Cosmic strins [62, 63, 64] are one of the topological defects that can arise in the early Universe when a symmetry is spontaneously broken by a complex scalar field. For the formation of CSs, the scalar field must have a vacuum manifold $\mathcal{M}$ so that the fundamental (or the first homotopy) group of $\mathcal{M}$ is non-trivial, i.e., $\pi_{1}(\mathcal{M})\neq I$. Generally, suppose a theory has a complex scalar $\phi$ with a quartic ‘Mexican hat’ potential, which has a degenerate circle of minima $|\phi|=\eta/\sqrt{2}$. In that case, the required condition for forming CSs can be satisfied. These CSs can be present as remnants of the symmetry breaking in the broken phase of the Universe.
@@
where $\rho_{GW}$ is the GW energy density and $\rho_{c}=(\sqrt{3}H_{0})^{2}/8\pi G$ is the critical energy density. We can decompose the total GW radiation from a loop into a sum of normal-mode oscillations with oscillation frequency $\tilde{f}=2k/\tilde{l}$, where $k\in\mathbb{Z}^{+}$. Here, $\tilde{l}\equiv\tilde{l}(\tilde{t})$ is the instantaneous loop-size at original GW-emission time, $\tilde{t}$. Taking the redshift into account, the observed frequencies today can be written as
@@ III.1 Gravitational Waves from global cosmic strings
[…]
@@
The GCS-mediated GW-spectrum has a vast frequency range, from $f_{0}\simeq 2/\alpha t_{0}\sim 3.6\times 10^{-16}$ Hz to a very large frequency depending on the symmetry breaking scale. For our purpose, we are interested in $f_{0}<f<f_{eq}$ (with $f_{eq}\sim 1.8\times 10^{-7}$ Hz) as PTA experiments like NANOGrav are sensitive to this range. The spectrum falls as $f^{-1/3}$ in this range and can be approximated by [67]
$h^{2}\Omega_{\text{GW}_{\text{GCS}}}(f)\simeq 2.9\times 10^{-12}{\left(\frac{\eta}{10^{15}\text{GeV}}\right)}^{4}{\left(\frac{f}{f_{eq}}\right)}^{-1/3}\,.$ (22)

### v2

@@ Abstract
We study the cosmology of a modified majoron model motivated by the need to protect a global $U(1)$ symmetry from gravity-induced hard explicit breaking (by $d\leq 4$ operators) at the Planck scale. The model extends the Standard Model by introducing a gauged $U(1)_{B-L}$ and an approximate global $U(1)$ symmetry, each spontaneously broken by a corresponding complex scalar singlet. This setup gives rise to a network of effectively global and local cosmic strings, whose stochastic gravitational wave signals can jointly account for the spectrum observed by the NANOGrav collaboration, particularly for majoron masses $m_{\chi}<10^{-23}$ eV. Although the fit is not as strong as that from supermassive black hole mergers, the model still provides an alternative explanation rooted in high-energy physics. The majoron mass induces an infrared cutoff in the gravitational wave spectrum, and we examine how this effect shapes the fit to the NANOGrav data. The model also generates light neutrino masses via the seesaw mechanism and avoids cosmological constraints from $\Delta N_{\text{eff}}$, CMB anisotropies, and isocurvature fluctuations. Although the majoron can contribute to dark matter through thermal, coherent oscillation, and string-induced production mechanisms, its relic abundance remains subdominant in the NANOGrav-compatible region. In contrast, the measured dark matter relic density is achievable at higher $m_{\chi}$, though at the cost of tension with cosmological bounds. If the NANOGrav fits are viewed as constraints, given their comparatively lower Bayes factors, they yield bounds that are significantly stronger than those imposed by the CMB and other cosmological data.
@@ I Introduction
The first direct detection of gravitational waves (GWs) by the LIGO–Virgo collaboration in 2015 [1] opened a new era in observational astronomy. In 2023, pulsar timing array (PTA) collaborations— the Parkes PTA [2], the European PTA [3], the Chinese PTA [4], and North American Nano-hertz Observatory for Gravitational Waves (NANOGrav) [5, 6]—reported evidence for a stochastic GW background in the nano-Hertz range. PTAs detect such signals by monitoring the arrival-time fluctuations of radio pulses from millisecond pulsars across the galaxy [7, 8, 9], with angular correlations following the Hellings–Downs curve [10] expected from spin-2 GWs. Among these, NANOGrav provides the most compelling evidence using 67 pulsars. While mergers of supermassive black hole binaries (SMBHBs) remain the leading astrophysical explanation [11, 12], no individual SMBHB mergers have been detected, motivating interest in cosmological sources [13, 14]. These include domain walls [15, 16, 17, 18], phase transitions [19, 20], scalar-induced GWs [21, 22, 23, 24], and cosmic strings [25, 26]. In this work, we focus on cosmic strings (CSs), which not only provide a plausible source of the signal but also offer a rare probe of high-scale physics, spanning $10^{12}$$10^{15}$ GeV (depending on local or global CS), well beyond the reach of current and proposed future colliders.
A recent work on global cosmic strings (GCSs) [27] shows that their GW spectrum, for symmetry-breaking scales around $\mathcal{O}(10^{15})$ GeV, can fall within the nano-Hertz range probed by PTAs. Intriguingly, the classic and simplest majoron model [28]—a well-known framework for neutrino mass generation—features a global $U(1)_{L}$ symmetry broken at $\mathcal{O}(10^{14}$$10^{15})$ GeV by a complex scalar $\phi$. This symmetry breaking has two key implications: (i) it generates GCSs that may emit GWs detectable by PTAs such as NANOGrav, and (ii) it induces high-scale Majorana masses for right-handed neutrinos (RHNs), enabling light neutrino mass generation via the Type-I seesaw mechanism [29, 30, 31, 32, 33, 34, 35]. The majoron is identified with the phase of the scalar field $\phi$. While the GCS-induced GW spectrum, characterised by a negative spectral tilt, provides a less favourable fit to the latest data than SMBHB mergers, it still offers a promising opportunity to explore and constrain the parameter space of the majoron model. Additionally, majoron models also have a wide range of astrophysical, cosmological, and laboratory implications [36].
@@
To address these issues, we adopt a more robust framework, which we call the modified majoron model, by gauging the $U(1)_{B-L}$ symmetry and introducing an additional complex scalar $\phi^{\prime}$ [39]. This extended model includes both a local $U(1)_{B-L}$ and an approximate global $U(1)$ symmetry. Both $\phi$ and $\phi^{\prime}$ are charged under $U(1)_{B-L}$, with only $\phi$ carrying the global charge. Assigning (BL)(B-L) charge of 2 to $\phi$ enables Majorana mass generation for RHNs. The (BL)(B-L) charge of $\phi^{\prime}$ can be assigned such that the leading $U(1)_{B-L}$-invariant operator that breaks the global $U(1)$ symmetry arises only at dimension $d\geq 5$. The phase of $\phi$ still constitutes the majoron in this model and is the potential DM candidate. While the primary focus of this work is on the cosmological implications of the modified majoron model, we also briefly compare the results with the same from the simplest majoron model.
@@
The extended symmetry structure of the modified majoron model gives rise to the formation of intricate topological defects in the early Universe. When the local $U(1)_{B-L}$ symmetry breaks, it generates LCSs, followed by the formation of GCSs from the spontaneous breaking of the global $U(1)$. As the Universe cools down to a temperature where the Hubble constant is comparable to the majoron mass, explicit symmetry breaking becomes relevant, leading to the formation of domain walls (DWs). Each of these topological defects can emit gravitational radiation. The string-wall network’s fate hinges on how many walls attach to each string. If stable DWs form, they can dominate the Universe’s energy density, contradicting the CMB observations. However, if only a single DW attaches to either type of string, the network becomes unstable and promptly collapses [39, 62, 63]. In this paper, we analyse scenarios where only a single DW is formed so that the GW spectrum is generated predominantly from CSs. [^1] In particular, we consider the scenario in which one DW is attached to one GCS. In this case, minimization of Majoron gradient energy around the local strings suggests that no wall can consistently attach to a local string (see section III.3 for details), rendering the system unstable and leading to the eventual collapse of the domain wall [63]. By contrast, when a single DW is attached to a local string, more than one wall can be attached to the associated global string. To the best of our knowledge, the collapse of this particular complex wall-string network has not been explicitly demonstrated in dedicated numerical simulations. Given this absence of direct numerical confirmation, we restrict our analysis to the case of a single domain wall attached to global strings only; see Table 1 for details. An interesting aspect explored in this work is the role of the majoron mass in setting an infrared (IR) cutoff on the GW spectrum, which in turn influences the fit to the NANOGrav data.
Recently, Ref. [64] constructed a model gauging an anomaly-free linear combination of baryon and lepton numbers to ensure proton stability. This model also features a gravity-induced, explicitly broken accidental $U(1)_{L}$ global symmetry and the associated majoron. The authors explore the majoron as a DM candidate for $m_{\chi}>10^{-5}$ eV and consider minimal thermal leptogenesis. Our fit to the NANOGrav data extends to this class of model as well. However, we focus on majoron masses below $10^{-15}$ eV, necessary for the string-wall network to survive long enough for the CS GW spectrum to reach below the nano-Hz frequency range.
@@ II.1 The simplest majoron model
$\mathcal{L}_{\text{new}}=\frac{i}{2}\overline{N_{I}}\gamma^{\mu}\partial_{\mu}N_{I}+\frac{1}{4}\lambda_{IJ}\overline{N_{I}}(1+\gamma_{5})N_{J}\phi+\frac{1}{4}\tilde{\lambda}_{IJ}\overline{\nu_{I}}(1+\gamma_{5})N_{J}\tilde{\Phi}-V_{1}(\phi)+\text{h.c.}$ (1)
where $\tilde{\Phi}$ is the dual of the SM Higgs doublet, and $N_{I},\nu_{J}$ are 4-component Majorana spinors (i.e $N^{c}_{I}=N_{I}$ and $\nu^{c}_{J}=\nu_{J}$). The indices $I,J=1,2,3$ are the generation indices. $V_{1}(\phi)$ is the tree-level potential of the singlet scalar $\phi$. In general, the Majorana-type Yukawa coupling matrix $\lambda$ can contain off-diagonal entries. However, for simplicity, we again assume the $\lambda$ matrix to be diagonal (hereafter, $\lambda_{II}$ will be simplified to $\lambda_{I}$). In this model, the lepton number is a conserved charge since all the SM lepton doublets and RHNs possess $U(1)_{L}$ charge of $L=1$ and $\phi$ carries L=2L=-2. Consequently, $\phi$ interacts with RHNs only.
@@ II.2 The modified majoron model
The reduced coupling $\tilde{g}_{d,\text{grav}}$ is exponentially suppressed by the combined scalar and gravitational Euclidean action in the region near the wormhole and is expected to be very small. For d=520d=5-20 and $\lambda=1$, $\tilde{g}_{d,\text{grav}}\sim 10^{-(d+2)}$. Given the structure of the effective Lagrangian in Eq. (4), it is unclear whether a single scalar field can suppress $d\leq 4$ effective operators and thus prevent explicit hard breaking of the global symmetry at the Planck scale.
To avoid this lack of clarity, we prefer a conservative approach and choose a model where a local gauge symmetry ensures that no $d\leq 4$ effective operators are induced by gravity. Such a model was constructed in Ref. [39] by gauging $U(1)_{B-L}$ and introducing an additional complex scalar field, $\phi^{\prime}$. The charge assignment of $\phi^{\prime}$ under (BL)(B-L) can be chosen such that the lowest-dimensional $U(1)_{B-L}$-invariant operator that breaks the global $U(1)$ symmetry appears only at $d\geq 5$.
In a nutshell, this model extends the SM gauge group by introducing a local $U(1)_{B-L}$ symmetry and an approximate global $U(1)$ symmetry. The particle content is extended by three RHNs and two complex scalars, $\phi$ and $\phi^{\prime}$, which are all singlets under the SM gauge groups. Here, both $\phi$ and $\phi^{\prime}$ are charged under the local $U(1)_{B-L}$ symmetry, and only $\phi$ is charged under the accidental global symmetry. If one assigns (BL)(B-L) charge of $q=2$ to $\phi$, it again couples with RHNs and generates Majorana masses for them. On the other hand, we keep $q^{\prime}$, the (BL)(B-L) charge of $\phi^{\prime}$, as general as possible (including fractional charges) to allow us to construct different models for a given $d$-dimensional soft global symmetry breaking operator.
The lowest dimensional quantum gravity induced global symmetry breaking, but the (BL)(B-L) invariant operator is
$\mathcal{L_{\text{grav}}}=g_{\text{grav}}\frac{{\phi}^{n}{(\phi^{\prime*})}^{n^{\prime}}}{M_{Pl}^{d-4}}+\text{h.c.}\,$, (6)
@@
$V_{1}(\phi,\phi^{\prime})=-\tilde{\mu}^{2}|\phi|^{2}+\lambda|\phi|^{4}-\tilde{\mu}^{\prime 2}|\phi^{\prime}|^{2}+\lambda^{\prime}|\phi^{\prime}|^{4}+\lambda_{\text{int}}|\phi|^{2}|\phi^{\prime}|^{2}+\Bigg(g_{\text{grav}}\frac{{\phi}^{n}{(\phi^{\prime*})}^{n^{\prime}}}{M_{Pl}^{d-4}}+\text{h.c.}\Bigg)\,$, (7)
@@
Once the $U(1)_{B-L}$ charge of $\phi$ is fixed at $q=2$ from the requirement of Majorana mass generation of RHNs, one is left with the freedom of only one parameter choice at their disposal to introduce a global symmetry breaking term in dimension $d$. One can thus choose to vary the parameter $n$, which can admit positive integer values up to d1d-1, and the values of $n^{\prime}$ and $q^{\prime}$ can be updated accordingly using the dimensional and gauge invariance conditions mentioned above. Such terms in the Lagrangian will inevitably give rise to an $M_{Pl}^{d-4}$-suppressed majoron mass [39]. For example, in dimension-5 for $q=2$, the (BL)(B-L) invariant but global symmetry breaking term can arise if $q^{\prime}$ takes values {1/2, 4/3, 3, 8}. In contrast, if we assign a fractional $U(1)_{B-L}$ charge to $\phi^{\prime}$, say $q^{\prime}=1/8$, then the lowest dimension at which the global symmetry can break explicitly is 17. More phenomenologically motivated benchmark points are shown in Table 5 in Section VI.
@@ III Cosmic Strings and Gravitational Waves
Cosmic strings [67, 68, 69] are one of the topological defects that can arise in the early Universe when a symmetry is spontaneously broken by a complex scalar field. For the formation of CSs, the scalar field must have a vacuum manifold $\mathcal{M}$ so that the fundamental (or the first homotopy) group of $\mathcal{M}$ is non-trivial, i.e., $\pi_{1}(\mathcal{M})\neq I$. A theory containing a complex scalar field $\phi$ with a quartic ‘Mexican hat’ potential, characterized by a degenerate circle of minima $|\phi|=\eta/\sqrt{2}$, satisfies the necessary condition for cosmic string formation. These CSs can be present as remnants of the symmetry breaking in the broken phase of the Universe.
@@
[…]
@@ III.1 Gravitational Waves from global cosmic strings
$\Omega_{\text{GW}}^{(k)}(f)\Bigg|_{\text{GCS}}=\frac{\mathcal{D}_{\alpha}}{\rho_{c}}\frac{2k}{f}\frac{F_{\alpha}}{\alpha}\int_{t_{F}}^{t_{0}}d\tilde{t}\frac{\Gamma^{(k)}G\mu^{2}}{\left(\alpha+\Gamma G\mu+\frac{\Gamma_{g}}{2\pi N}\right)}\frac{\varepsilon_{\text{loop}}\left(t_{i}^{(k)}\right)}{{t_{i}^{(k)}}^{4}}\theta(\tilde{l})\theta(\tilde{t}-t_{i}){\left(\frac{a(\tilde{t})}{a(t_{0})}\right)}^{5}{\left(\frac{a(t_{i}^{(k)})}{a(\tilde{t})}\right)}^{3}\,$, (21)
@@
The GCS-mediated GW-spectrum has a vast frequency range, from $f_{0}\simeq 2/\alpha t_{0}\sim 3.6\times 10^{-16}$ Hz to a very large frequency depending on the symmetry breaking scale. For our purpose, we are interested in $f_{0}<f<f_{th}$ (with $f_{th}\sim 1.8\times 10^{-7}$ Hz) as PTA experiments like NANOGrav are sensitive to this range. The spectrum falls as $f^{-1/3}$ in this range and can be approximated by [72]
$h^{2}\Omega_{\text{GW}_{\text{GCS}}}(f)\simeq 2.9\times 10^{-12}{\left(\frac{\eta}{10^{15}\text{GeV}}\right)}^{4}{\left(\frac{f}{f_{th}}\right)}^{-1/3}\,.$ (22)